Text Size

February 20, 2010
ISS On-Orbit Status 02/20/10

All ISS systems continue to function nominally, except those noted previously or below. Sleep shift in effect: The ISS crew’s workday began last evening at 3:14pm and ended this morning at 6:44 EST. Currently, an extended sleep period (6:44am-6:00am) returns the wake/sleep cycle to “normal”. Saturday – Undock Day.

ISS and STS-130/Endeavour are flying in separate orbits again. [After hatch closure yesterday morning at 3:10am EST, Endeavour undocked last evening at 7:54pm from PMA-2 (Pressurized Mating Adapter 2). After separation, PLT Terry Virts completed a 270-deg station flyaround, and the Orbiter crew obtained photo/video imagery of the ISS. Final sep burn was at 9:37pm, posigrade.]
  • Endeavour is expected to return to Earth tomorrow (Sunday) after its 14-day mission. Landing opportunities at KSC are at 10:16pm & 11:51pm EST. There are additional opportunities at 1:20am & 2:55am Monday at EAFB (Edwards Air Force Base, Calif.), a backup landing site.

At wake-up last night, FE-4 Kotov did the regular daily “early-morning” check of the aerosol filters at the Russian Elektron O2 generator which Suraev had installed on 10/19 in gaps between the BZh Liquid Unit and the oxygen outlet pipe (filter FA-K) plus hydrogen outlet pipe (filter FA-V). [FE-4 again inspected the filters this morning before bedtime, currently a daily requirement per plan, with photographs to be taken if the filter packing is discolored.]

Crew activities before and during undocking included –
  • Setting up the departure configuration at the PMA-2 and standing by at the Lab RWS (Robotics Workstation) A31p laptop with a stopwatch to monitor the proper performance of automatic undocking software for the PMA-2 departure under Russian thruster attitude control and to immediately mode ISS to attitude control if undocking software fails (Williams). [The procedure provides for the crewmember to take over the automatic operational attitude control sequence manually if the software does not resume control after the period of free drift a few minutes after physical separation. Free drift is employed to prevent a conflict between the control systems of the two vehicles (ISS & Shuttle) and to “limp” (unload) the docking mechanisms],
  • Closing the protective window shutters on Kibo, Lab and Cupola (Williams), and
  • Using the NIKON digital camera and PD-100 camcorder to document the undocking, backing away & separation of the Endeavour (Noguchi).

FE-1 Suraev verified proper operation of the running Russian TEKh-15/DAKON-M IZGIB (“Bend”) experiment in the SM (Service Module) for taking structural dynamics data during the Shuttle undocking. Afterwards, Maxim downlinked the measurement data to the ground. [IZGIB has the objective to help update mathematical models of the ISS gravitation environment, using accelerometers of the Russian SBI Onboard Measurement System, the GIVUS high-accuracy angular rate vector gyrometer of the SUDN Motion Control & Navigation System and other accelerometers for unattended measurement of micro-accelerations at science hardware accommodation locations - (1) in operation of onboard equipment having rotating parts (gyrodynes, fans), (2) when establishing and keeping various ISS attitude modes, and (3) when performing crew egresses into space and physical exercises.]

After the undocking –
  • FE-4 Kotov restored the RS/USOS STTS communication system to its nominal post-undocking setup;
  • CDR deconfigured the BPSMU (Battery Powered Speaker Microphone Unit) and its long single drag-through line, used during the docked phase, stowing the equipment;
  • Jeff then deconfigured the C&T (Command & Tracking) video set up, uninstalling the video cap in Node-2 which enabled reception of video from the Endeavour with the Orbiter docked;
  • Williams also returned PEPs (Portable Emergency Provisions) back to their nominal stage positions, i.e., PBA (Portable Breathing Apparatus) #1019 from the A/L (Airlock) to COL (Columbus Orbital Laboratory) endcone & PBA #1028 from A/L to Kibo JPM (JEM Pressurized Module); and
  • Jeff depressurized the PMA-2 to prevent humidity condensation and pressure fluctuations, followed by leak checking for the standard one hour. [Afterwards, the necessary testing equipment was torn down, to be used next week for PMA-3.]

At ~1:09am this morning, the crew underwent the standard New Module Emergency OBT (Onboard Training), for the new additions Node-3 and Cupola. [Purpose of the one-hour drill was to familiarize the station crew with the location of Node-3 and Cupola hardware used in emergencies, and familiarization with the changes in the emergency response resulting from the addition of Node-3 and Cupola to the ISS configuration. The OBT included an inspection of all hatches and the ability to rapidly clear hatchways in emergencies, with detailed attention to cables and all other obstructions at hatchways, then to verify integrity of all hatch rubber seals and availability of foot restraints. Equipment checked for function and location included PBA (2), EKTK (Extension Hose/Tee Kit), PFE (Portable Fire Extinguisher), PBA oxygen ports (2), C&W (Caution & Warning) Panel, and ATUs (Audio Terminal Units).]

Suraev & Kotov had several hours set aside to stow excessed equipment and trash in Progress 36P, to be jettisoned later this summer. [The moves were logged in the IMS (Inventory Management System.]

Maxim retrieved the BIO-2 Biorisk-MSN container, brought aboard during the Russian EVA-24, from stowage, took it out of its container and photographed it from different angles, to update its documented configuration after the spacewalk. [Biorisk-MSN will be returned on Soyuz TMA-16/20S on 3/18.]

The FE-1 also completed the regular maintenance of the Russian POTOK-150MK (150 micron) air purification subsystems in the SM & FGB, cleaning the pre-filters with a vacuum cleaner with narrow nozzle attachment and later restarting POTOK in automatic mode.

FE-4 Kotov took documentary photography of the area behind the panel for the installation of the BZU onboard memory device.

Oleg also performed the periodic accuracy check on MV vacuum pressure gauges in the RS (6).

Also in the SM, Suraev conducted periodic maintenance on the deactivated Russian IK0501 GA (Gas Analyzer) of the SOGS Pressure Control & Atmospheric Monitoring System behind panel 449 by replacing its CO2 filter assembly (BF) with a new unit from FGB stowage (done last: 1/12/10), then reactivating the unit for a calibration check with its IG-3 oxygen (O2) sensor unit, supported by tagup with ground specialists. [IK0501 is an automated system for measuring CO2, O2, and H2O in the air, as well as the flow rate of the gas being analyzed.]

In the Japanese JPM, Soichi Noguchi installed a CO2 valve on the CGSE (Common Gas Supply Equipment) to supply CO2 to the MyoLab experiment during the 19A mission.

Also in the Kibo module, Soichi took the periodic photography of the Dewey’s Forest science payload, which requires periodic watering for the cultivation of its PUs (Plant Units). [Dewey’s Forest, one of the Japanese educational payloads, is intended to show how gravity controls the laws of nature and influences our ways of thinking. The project is “a catalyst to rediscover our relationship with plants on the ground and the age-old history of our gardens.”]

The crewmembers worked out with their regular 2-hr physical exercise on the CEVIS cycle ergometer (CDR), TVIS treadmill (FE-1, FE-4), ARED advanced resistive device (CDR, FE-5, FE-6), T2/COLBERT advanced treadmill (FE-5, FE-6), and VELO bike ergometer with bungee cord load trainer (FE-1, FE-4).\

  • Sleep shifting started with the Progress docking on 2/4 and is now returning to Normal. Revised Wake/Sleep table:

3:14pm (2/19)

Weekly Science Update (Expedition Twenty-Two -- Week 12)

3-D SPACE: No report.

AgCam (Agricultural Camera): No report.

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Complete.


APEX (Advanced Plant Experiments on Orbit) -Cambium: No report.

APEX-TAGES (Transgenic Arabidopsis Gene Expression System): The first of three harvests will happen next week.

BCAT-4/5 (Binary Colloidal Alloy Test 4/5): No report.

BIOLAB (ESA): No report.

BIORHYTHMS (JAXA, Biological Rhythms): No report.

BISE (CSA, Bodies in the Space Environment): No report.

BISPHOSPHONATES: “Jeff: thanks for completing your pill ingestion. Your next session is scheduled for 2/21. Soichi: thanks for completing your pill ingestion. Your next session is scheduled for 2/21.”

CARD (Long Term Microgravity: Model for Investigating Mechanisms of Heart Disease, ESA): No report.

CARDIOCOG-2: Complete.

CB (JAXA Clean Bench): No report.

CBEF (JAXA Cell Biology Experiment Facility)/SPACE SEED: “Soichi, thank you for successfully fixing the fan in the 1G compartment and installing a new Air control Fan. The repair was checked and confirmed that this configuration can support CBEF temperature control. Good job.”

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): No report.

CERISE (JAXA): Frozen Samples moved to GLACIER. Sample will be returned by 20A.

CETSOL (ESA): 2 processed sample cartridge assemblies (SCAs) will be downloaded on 20A.

CFE (Capillary Flow Experiment): Reserve.

CSI-3/CGBA-5 (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus 5): Ongoing.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CIR (Combustion Integrated Rack), MDCA/Flex: No report.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): Complete.

Commercial Photo (JAXA): No report.
CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): Complete.

DECLIC (Device for the Study of Critical Liquids & Crystallization, CNES/NASA): No report.

DomeGene (JAXA): Complete.

DOSIS (ESA): No report.

EarthKAM (Earth Knowledge Acquired by Middle School Students): No report.

EDR (European Drawer Rack, ESA): No report.

ELITE-S2 (Elaboratore Immagini Televisive - Space 2): Planned.

EMCS (European Modular Cultivation System): No report.

ENose (Electronic Nose): No report.

EPM (European Physiology Module): No report.

EPO (Educational Payload Operations, NASA): No report.

EPO LES-2 (ESA): No report.

EPO 3-min Video (JAXA): No report.

EPO J-Astro Report (JAXA): Ongoing.

EPO Dewey’s Forest (JAXA): Watering in progress.

EPO Space Clothes (JAXA): Complete.

EPO Hiten (Dance, JAXA): No report.

EPO Moon Score (JAXA): No report.

EPO Try Zero-G (JAXA): No report.

EPO Kibo Kids Tour (JAXA): Complete.

EPO Paper Craft (Origami, JAXA): No report.

EPO Poem (JAXA): Photo-taking activity will be a Voluntary Science candidate.

EPO Spiral Top (JAXA): No report.

ETD (Eye Tracking Device): Completed.

FACET (JAXA): No report.

FIR/LMM/CVB (Fluids Integrated Rack / Light Microscopy Module / Constrained Vapor Bubble): No report.

FWED (Flywheel Exercise Device, ESA): No report.

FOAM STABILITY (ESA): Tapes were returned on ULF3.

FOCUS: “Jeff, both raw and jpeg images have been retrieved from the restricted area of the Digital Imagery Management System (DIMS), and transferred to the FOCUS science team for detailed analysis. First feedback is very positive.”

FSL (Fluid Science Laboratory, ESA): No report.

GEOFLOW: No report.

HAIR (JAXA): Soichi collected hair sample from TJ on 1/19. “Thank you for bearing the pain.”

HDTV System (JAXA): Was delivered by HTV1.

Holter ECG (JAXA): No report.

HQPC (JAXA): Was delivered by 34P.

HREP (HICO/Hyperspectral Imager for the Coastal Ocean & RAIDS/Remote Atmospheric & Ionospheric Detection System/JAXA): Due to power constraints, HREP was placed in low power mode on 2/12. Survival power and avionics remain on.


ICV (Integrated Cardiovascular): “Jeff, with the additional Actiwatch and Holter card downloads, we have been able to verify that all of your FD75 and FD135 data is now on the ground. The PIs have looked through the Cardiopres data for the FD75 session and, while it was not possible to record for 24 hours, the data that was recorded is usable.”

IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Complete.


InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): No report.

IRIS (Image Reversal in Space, CSA): No report.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): No report.

Marangoni UVP (JAXA): No report.

MAXI (Monitor of All-sky X-ray Image, JAXA): Continuing observation operation.

Marangoni UVP: Going well. Oscillation flow is under observation.

MEIS (Marangoni Experiment for ISS) in JAXA FPEF (Fluid Physics Experiment Facility): No report.

MDCA/Flex: See under CIR.

Microbe-1 (JAXA): No report.

Micro-G Clay (JAXA EPO): Complete.

MMA (JAXA/Microgravity Measurement Apparatus): No report.

MISSE7 (Materials ISS Experiment): To prevent overheating during the ISS rotation and shuttle docking, MISSE7 PEC 7b and 7a nadir were shut down on 2/9. The PECs are planned to be returned to normal operations on 47/20:00.

MSG-SAME (Microgravity Science Glovebox): Complete.

MSL (Materials Science Laboratory): No report.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC-1 “Pirs”.

MULTIGEN-1: Completed.

MYCO 1 (JAXA): Sample Collection and MELFI stow are complete. Sample will return on 20A using DCB (Double Cold Bag).

NEUROSPAT (ESA/Study of Spatial Cognition, Novelty Processing and Sensorimotor Integration): No report.

NOA-1/-2 (Nitric Oxide Analyzer, ESA): Complete.


PADLES (JAXA, Area PADLES 3; Passive Area Dosimeter for Lifescience Experiment in Space): No report.

PASSAGES (JAXA): No report.

PCDF-PU (Protein Crystallization Diagnostic Facility - Process Unit): No report.

PCG (JAXA, Protein Crystal Growth): No report.

PCRF (Protein Crystallization Research Facility) Reconfiguration (JAXA): Complete.

PMDIS (Perceptual Motor Deficits in Space): Complete.


RadGene & LOH (JAXA): Complete.

RadSilk (JAXA): Samples were returned to ground on ULF3.

RST/Reaction Self Test (Psychomotor Vigilance Self Test on the ISS): “Jeff, TJ, and Soichi: thank you for your continued participation in Reaction Self Test!”

SAIBO Rack (JAXA): No report.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): No report.

SAMPLE: Complete.

SCOF (Solution Crystallization Observation Facility, JAXA): No report.

SEDA-AP (Space Environment Data Acquisition Equipment-Attached Payload, JAXA): Continuing observation operations.

SHERE (Shear History Extensional Rheology Experiment): Complete.

SLAMMD (Space Linear Acceleration Mass Measurement Device): No report.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): “Jeff and TJ: Thanks for completing another week of Sleep logging.”

SMILES (JAXA): Continuing observations.

SODI/IVIDIL (Selectable Optical Diagnostics Instrument/Influence of Vibration on Diffusion in. Liquids, ESA): “Thank you Soichi for completing the last steps of the SODI setup. With the installation of the DSC Cell Array#1 and the first Flash Disk, we are ready to begin the science runs.”

SOLAR (Solar Monitoring Observatory): Sun observation window #25 started on 2/9. The tracking windows on each orbit have been shorter than usual due to the attitude of the ISS. This on-going Sun observation window is expected to end by 2/20.

SOLO (Sodium Loading in Microgravity): No report.

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): No report.

SPICE (Smoke Point In Co-flow Experiment): No report.:

SPINAL (Spinal Elongation): “Jeff, TJ, and Soichi: Thank you for working together to complete the Spinal Elongation activities. We have received the photos and are in the process of analyzing. Also, thank you for going above and beyond by taking pictures of Standing Height, it is great to have insight into how it is performed. We will receive the Seated Height and Standing Height logs after STS 130/20A landing. We look forward to the analysis of all the data.”

SWAB (Characterization of Microorganisms & Allergens in Spacecraft): No report.

THERMOLAB (ESA): No report.

TRAC (Test of Reaction & Adaptation Capabilities): Planned.


VASCULAR (CSA): “Jeff, thanks for completing your blood draw. The PI team will be in Russia to welcome you back and complete post-flight BDC with you.”
VO2max (NASA): No report.

VLE (Video Lessons ESA): No report.

WAICO #1/#2 (Waving and Coiling of Arabidopsis Roots at Different g-levels): No report.

YEAST B (ESA): No report.

CEO (Crew Earth Observations): No report.

No CEO photo targets uplinked for today.

Significant Events Ahead (all dates Eastern Time and subject to change):
02/20/10 – ISS Stage Reboost (~3:00pm)
02/21/10 -- STS-130/Endeavour/20A KSC deorbit burn (9:13pm or 10:49pm)
02/21/10 -- STS-130/Endeavour/20A KSC landing (10:16pm or 11:51pm)
03/18/10 -- Soyuz TMA-16/20S undock/landing
--------------Three-crew operations-------------
03/18/10 -- STS-131/Discovery/19A – MPLM(P), LMC (launch ~1:30pm EST)
04/02/10 -- Soyuz TMA-18/22S launch – Skvortsov (CDR-24)/Caldwell/Kornienko
04/04/10 -- Soyuz TMA-18/22S docking
--------------Six-crew operations-----------------
04/27/10 -- Progress M-03M/35P undock
04/28/10 -- Progress M-05M/37P launch
04/30/10 -- Progress M-05M/37P docking
05/14/10 -- STS-132/Atlantis/ULF4 – ICC-VLD, MRM-1 (~2:00pm EST)
05/10/10 -- Progress M-04M/36P undock
05/31/10 -- Soyuz TMA-17/21S undock/landing
--------------Three-crew operations-------------
06/14/10 -- Soyuz TMA-19/23S launch – Wheelock (CDR-25)/Walker/Yurchikhin
06/16/10 -- Soyuz TMA-19/23S docking
--------------Six-crew operations-----------------
07/xx/10 -- US EVA-15
07/xx/10 -- Russian EVA-25
06/28/10 -- Progress M-06M/38P launch
07/02/10 -- Progress M-06M/38P docking
07/26/10 -- Progress M-05M/37P undock
07/27/10 -- Progress M-07M/39P launch
07/29/10 -- Progress M-07M/39P docking
07/29/10 -- STS-134/Endeavour (ULF6 – ELC3, AMS-02) (~7:30am EST)
08/30/10 -- Progress M-06M/38P undock
08/31/10 -- Progress M-08M/40P launch
09/02/10 -- Progress M-08M/40P docking
09/15/10 -- Soyuz TMA-18/22S undock/landing
09/16/10 -- STS-133/Discovery (ULF5 – ELC4, PMM) (~12:01pm EST)
09/18/10 -- STS-133/Discovery (ULF5 – ELC4, PMM) docking
09/22/10 -- STS-133/Discovery (ULF5 – ELC4, PMM) undock
09/30/10 -- Soyuz TMA-20/24S launch – Kelly (CDR-26)/Kaleri/Skripochka
10/xx/10 -- Russian EVA-26
10/26/10 -- Progress M-07M/39P undock
10/27/10 -- Progress M-09M/41P launch
10/29/10 -- Progress M-09M/41P docking
11/15/10 -- Soyuz TMA-19/23S undock/landing
11/30/10 -- Soyuz TMA-21/25S launch – Kondratyev (CDR-27)/Coleman/Nespoli
12/15/10 -- Progress M-08M/40P undock
02/08/11 -- Progress M-09M/41P undock
02/09/11 -- Progress M-10M/42P launch
02/11/11 -- Progress M-10M/42P docking
03/30/11 -- Soyuz TMA-22/26S launch – A. Borisienko (CDR-28)/R, Garan/A.Samokutayev
xx/xx/11 -- Progress M-11M/43P launch
05/30/11 -- Soyuz TMA-23/27S launch – M. Fossum (CDR-29)/S. Furukawa/S. Volkov
12/??/11 -- 3R Multipurpose Laboratory Module (MLM) w/ERA – on Proton.